OpenStudy (anonymous):

What is the vertex for y=3/4x^2?

OpenStudy (nottim):

Are you allowed drawing the graph of the equation?

OpenStudy (anonymous):

yes

OpenStudy (nottim):

Have you done so yet? I can't remember how to do this....

OpenStudy (mathmate):

y=3/4x^2 = (3/4)(x-0)+0 or the vertex is at (0,0)

OpenStudy (across):

Develop an intuition for the problem: you can tell the vertex is at (0, 0) in less than a second.

OpenStudy (mathmate):



OpenStudy (mathmate):

Yes, (3/4)x^2 is a multiple of y=x^2, so the vertex does not change.

OpenStudy (anonymous):

so vertex would be y= ax^ vertex =a ?

OpenStudy (anonymous):

y= ax^2

OpenStudy (mathmate):

The "a" will magnify the height of the parabola, but the vertex will remain at (0,0). Think of the vertex equation: y=a(x-h)+k If h and k are zero, then other values of a will not change the vertex, which is at (h,k).

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

so is just like the vertex of a triangle kind of :P